High frequency sound can be used to produce  standing wave vibrations in a wine glass. The glass shatters when the wave is  observed to have four nodes and 4 anti nodes equally spaced around the 20.0 cm long  glass.   If transverse waves moves around the glass at 900m/s, with what frequency will an opera singer to shatter the glass?




1
Expert's answer
2021-02-22T10:27:44-0500
"\lambda=\dfrac{L_{circumference}}{2}=\dfrac{0.2\ m}{2}=0.1\ m,""v=f\lambda,""f=\dfrac{v}{\lambda}=\dfrac{900\ \dfrac{m}{s}}{0.1\ m}=9000\ Hz."